# Gaussian process regression with Sliced Wasserstein Weisfeiler-Lehman graph kernels

### 06 Feb 2024 v1

@@ 2 RELATED WORK
Recently, many approaches focused on solving the graph-alignment problem using optimal transport and trying to find a matching between well-chosen distributions. The Fused-Gromov distance proposed by Vayer et al., (2018) makes it possible to separate the information provided by the graph structure and its attributes with a trade-off between Wasserstein and Gromov-Wasserstein distances. The recent approach of Kaloga et al., (2022) uses embeddings obtained with Simple Graph Convolutional Network (SGCN) (Wu et al.,, 2019), a simplified version of GCN reducing the number of trainable parameters. They also replace the Wasserstein distance by the Restricted Projected Wasserstein that uses projections to determine the optimal transport plans and their costs. Other approaches propose to use the Linear Optimal Transport framework by computing transport displacements to a reference template (Kolouri et al., (2020)) or to regularize the Wasserstein distance using both local and global structural information of the graph (Wijesinghe et al., (2021)). However, all approaches using Wasserstein or Gromov-Wasserstein distances suffer from scalability issues and from the impossibility to obtain a positive definite kernel.
Graphs aside, many approaches tackle the problem of distribution regression, where a function over spaces of probabilities is learnt (Smola et al.,, 2007; Póczos et al.,, 2013). There are some positive definite kernels using optimal transport with possible entropic regulariation Bachoc et al., (2020, 2023) but they require the use of a reference measure. Alternatively, the sliced Wasserstein distance can be used to define positive definite distance substitution kernels. It was initially proposed by Kolouri et al., (2016) in order to build kernel functions between absolutely continuous distributions with respect to the Lebesgue measure and was extented to empirical distributions by Meunier et al., (2022). Our work uses the results of Meunier et al., (2022) on the SW substitution kernel but differs in the SW approximation.
Finally, note that the notion of kernels between graphs is sometimes extended to fingerprint kernels for molecular data using for instance the Tanimoto/Jaccard kernel (Griffiths et al.,, 2022). However, such molecular kernels can only be used for this type of data and require a specific vector representation of the graphs, which is usually a complex problem.
@@ 4 SWWL GRAPH KERNEL
The proofs are given in the Supplementary Material (the positive definiteness result relies on the work on Meunier et al., (2022)). We now detail the construction of the feature map $\bm{\phi}$ involved in Eq. (5).
@@ 5.1 Classification
The proposed SWWL kernel is compared with 5 existing graph kernels, namely, WWL (Togninalli et al.,, 2019), Fused Gromov Wasserstein (FGW) (Vayer et al.,, 2018), Restricted Projected Wasserstein (RPW) (Kaloga et al.,, 2022), Propagation (PK) (Neumann et al.,, 2016) and Graph Hopper (GH) (Feragen et al.,, 2013). Five datasets are taken from the TUD benchmark (Morris et al.,, 2020): BZR, COX2, PROTEINS, ENZYMES, Cuneiform. We only keep continuous attributes for the nodes and ignore categorical information except for Cuneiform where we use the same setup as proposed by Kaloga et al., (2022) with fused attributes.
@@ 5.2 Regression
Rotor37 dataset. The NASA rotor 37 case serves as a prominent example of a transonic axial-flow compressor rotor widely employed in computational fluid dynamics research. This dataset is made of 3D compressible steady-state Reynold-Averaged Navier-Stokes (RANS) simulations that model external flows (Ameri,, 2009). The inputs of the simulations are given by the finite element mesh of a 3D compressor blade (graphs with $\sim 30000$ nodes), and two scalar physical parameters corresponding to the rotational speed and the input pressure. The scalar output of the problem is the isentropic efficiency, which is computed by solving the boundary value problem with the finite volume method using the elsA software (Cambier et al.,, 2011).
@@
AirfRANS dataset. This dataset consists of RANS simulations over 2D NACA airfoils with geometric variabilities in a subsonic flight regime (Bonnet et al.,, 2022). The inputs of the problem are given by the finite element mesh of the domain surrounding the airfoil (graphs with more than $10^{5}$ nodes), and two scalar inputs, the angle of attack and the inlet velocity. The scalar output of the problem is the drag coefficient. More details about the distributions of the training and test sets can be found in the work of Bonnet et al., (2022).
@@ Acknowledgements
This work was supported by the French National Research Agency (ANR) through the SAMOURAI (Simulation Analytics and Metamodel-based solutions for Optimization, Uncertainty and Reliability AnalysIs) project under grant ANR20-CE46-0013.
@@
Appendix A Optimal transport: definitions and properties
@@
Appendix B Proofs
B.1 Positive definiteness of SWWL
Definition 7 (Hilbertian (pseudo)-distance).
((Hein and Bousquet,, 2005)) A pseudo-distance $d$ defined on a set $X$ is said to be Hilbertian if there exists a Hilbert space $\mathcal{H}$ and a feature map $\phi:X\rightarrow\mathcal{H}$ such that for any pair $x,x^{\prime}$ in $X$, $d(x,x^{\prime})=||\phi(x)-\phi(x^{\prime})||_{\mathcal{H}}$.
Pseudo means that $d$ might not satisfy $d(x,x^{\prime})=0$ implies $x=x^{\prime}$. All usual kernels substituted with distances are positive definite if and only if the distance is Hilbertian as the following property shows.
Property 4.
Meunier et al., (2022) Introducing $\langle x,x^{\prime}\rangle^{x_{0}}_{d}=\langle\phi(x)-\phi(x_{0}),\phi(x^{\prime})-\phi(x_{0})\rangle_{\mathcal{H}}$, the following statements are equivalent:
  • $d$ is a Hilbertian pseudo-distance.
Property 5.
(Peyré et al., (2019)) $\sqrt{W_{1}}$ and $W_{2}$ are not Hilbertian distances in $\mathbb{R}^{s}$ when $s\geq 2$.
Property 6.
(Meunier et al., (2022)) $\sqrt{SW_{1}}$ and $SW_{2}$ are Hilbertian distances in $\mathbb{R}^{s}$ for all $s\geq 2$.
We can now prove the positive definiteness of SWWL.
Proof.
We recall the form of the estimated sliced Wasserstein distance between two distributions $\mu$ and $\nu$ with $P$ projections and $Q$ quantiles: $\widehat{SW}_{r,P,Q}(\mu,\nu)=||\phi(\mu)-\phi(\nu)||_{r}$ where ||.$||.||_{r}$ is the $r$-norm in $\mathbb{R}^{PQ}$ and $\bm{\phi}_{p+P(q-1)}(\mu)=(PQ)^{-1/r}F^{-1}_{\mu_{\theta_{p}}}(t_{q})$ for $p=1,\dots,P$ and $q=1,\dots,Q$. $\widehat{SW}_{r,P,Q}$ is a Hilbertian pseudo-distance 7 as we built an explicit feature map $\phi$ to $\mathcal{H}=\mathbb{R}^{PQ}$.
Finally, this property still holds if we restrict the space of measures to be $\{\mu_{G}=|V|^{-1}\sum\limits_{u\in V}\delta_{\mathbf{E}^{G}_{u}},G\in\mathcal{G}\}$, i.e. the empirical distributions associated with the continuous WL embeddings of graphs in $\mathcal{G}$.
B.2 Complexity of SWWL
In this section we give a proof of Property 2 (Complexity of SWWL).
Proof.
In comparison, using the Wasserstein distance would lead to a $O(N^{2}n^{3}log(n))$ time complexity and a $O(n^{2})$ space complexity to build the Gram matrix, which can be prohibitive when the number of nodes is large.
@@ Appendix C Details about the compared methods and their hyperparameters
In this section we give more details on hyperparameters of reference methods for classification and regression tasks presented in the experimental section. We also explain the choice of the compared methods for regression. In particular, we emphasize the difficulties of applying other approaches to the large meshes of interest, hence we also consider coarsened meshes which make it possible to compare the different methods. Please note that we used indefinite kernels for the kernels methods (SVM and Gaussian process regression), even if the the existence of a reproducing kernel Hilbert space in this case isn’t true. Some work try to extend to the indefinite case in Krein spaces, but it remains an open question (see for example the work of Liu et al., (2020) for kernel ridge regression). We nevertheless test several graph kernels regardless of the positive definiteness as they are commonly used and give good results in practice.
@@
SWWL (ours) $\mathcal{O}(NH\delta n$ + $NPn(log(n)+H)+N^{2}PQ)$ $P$: projections, $Q$: quantiles, $H$: WL iterations
WWL $\mathcal{O}(NH\delta n+HN^{2}n^{3}log(n))$ $H$: WL iterations
@@
RPW $\mathcal{O}(NHn^{2}+N^{2}Pnlog(n))$ $H$: layers, $P$: last layer size
@@
Two major problems arise when comparing to WWL. The time to calculate a Wasserstein distance between embeddings of size 30000 is approximately 80 seconds, so more than 400 days are needed to compute the pairwise distances of 1000 graphs. The second issue is memory usage. Allocating matrices of size 30000 $\times$ 30000 for the transport costs of the Wasserstein distance is only possible with a minimum amount of space and severely limits parallelization. We still include a comparison with WWL on coarsened datasets with 100 parallel jobs as it is the method on which SWWL is based. Remark that we do not use entropic regularization (Cuturi, (2013)) that would reduce the complexity but has the (not negligible) cost of an extra hyperparameter to tune in an outer loop.
@@
The datasets Rotor37,Tensile2d, and AirfRANS can be found according to the link given in the main paper. Please note that the datasets Rotor37 that we used in the exeperiments and the one that is publiquely available use different type of finite elements (quads/triangles) and the conversion from quads to triangles yields possibly different graphs. Scores may vary slightly with this public version.
@@ C.3 Robust Gaussian process regression
In the robust Gaussian process framework, the marginal likelihood is augmented by a prior for the range parameters $\pi(\theta,\sigma^{2},\bm{\gamma})\propto\frac{\pi(\mathbf{\gamma)}}{\sigma^{2}}$. We use the prior $\pi^{JR}(\bm{\gamma})=\left(\sum_{l=1}^{L}C_{l}\gamma_{l}^{-1}\right)^{a}\exp\left(-b\sum_{l=1}^{L}C_{l}\gamma_{l}^{-1}\right)$, $a=0.2$, $b=N^{-1/L}(a+L)$, and $C_{l}$ equal to the mean of $|x_{il}-x_{il}|$ for $1\leq i,j\leq N,i\neq j$. Remark that in the case where $x_{il}$ is not a scalar but rather a structured object, we replace the absolute value by the user-defined distance or pseudo-distance that is substituted in the kernel ((in our paper we use the estimated sliced Wasserstein distance).
@@
$(\hat{\gamma_{1}},\ldots,\hat{\gamma_{L}})=\operatorname*{argmax}\limits_{\gamma_{1},\ldots,\gamma_{L}}(\mathcal{L}(\gamma|\mathbf{X},\mathbf{y})\pi^{JR}(\bm{\gamma})).$ (16)
We now want to predict unknown targets $\mathbf{f}_{*}:=(f(\mathbf{x}^{*}_{i}))_{i=1}^{N^{*}}$ at $N^{*}$ new locations $\mathbf{X}^{*}=(\mathbf{x}^{*}_{i})_{i=1}^{N^{*}}$. With the use of the estimated range parameters, the predictive distribution of $\mathbf{f}^{*}$ given $\mathbf{X}$, $\mathbf{y}$ and $\bm{\gamma}$ is a Student’s $t$-distribution with $N-1$ degrees of freedom
@@
with $\mathbf{h}_{*}$ being the $N^{*}\times 1$ vector composed of ones, $\hat{\theta}=(\mathbf{h}^{T}\mathbf{R}^{-1}\mathbf{h})^{-1}\mathbf{h}^{T}\mathbf{R}^{-1}\mathbf{y}$ being the least squares estimator for $\theta$, and $\mathbf{R}_{*}$, $\mathbf{R}_{**}$ being the covariance matrices evaluated respectively between train inputs and test/train inputs.
@@
We add the times required to train the Gaussian process regression with RobustGaSP in R given precomputed distance matrices. For Rotor37 (1000 train inputs), the training takes 220 seconds, for Tensile2d (500 train inputs), for AirfRANS (800 train inputs) the training takes 135 seconds. This time is similar for coarsened meshes as the shape of distance matrices is the same.
@@ D.1 Anisotropic variant of SWWL
$k_{ASWWL}(G,G^{\prime})=\prod_{h=0}^{H}\exp\left(-\gamma_{h}\widehat{SW}^{2}_{2,P,Q}(\mu^{(h)}_{G},\mu^{(h)}_{G^{\prime}})\right),$ (21)

### 11 Mar 2024 v2

@@ 2 RELATED WORK
Recently, many approaches focused on solving the graph-alignment problem using optimal transport and trying to find a matching between well-chosen distributions. The Fused-Gromov distance proposed by Titouan et al., (2019) makes it possible to separate the information provided by the graph structure and its attributes with a trade-off between Wasserstein and Gromov-Wasserstein distances. The recent approach of Kaloga et al., (2022) uses embeddings obtained with Simple Graph Convolutional Network (SGCN) (Wu et al.,, 2019), a simplified version of GCN reducing the number of trainable parameters. They also replace the Wasserstein distance by the Restricted Projected Wasserstein that uses projections to determine the optimal transport plans and their costs. Other approaches propose to use the Linear Optimal Transport framework by computing transport displacements to a reference template (Kolouri et al., (2021)) or to regularize the Wasserstein distance using both local and global structural information of the graph (Wijesinghe et al., (2021)). However, all approaches using Wasserstein or Gromov-Wasserstein distances suffer from scalability issues and from the impossibility to obtain a positive definite kernel.
Graphs aside, many approaches tackle the problem of distribution regression, where a function over spaces of probabilities is learnt (Smola et al.,, 2007; Póczos et al.,, 2013). There are some positive definite kernels using optimal transport with possible entropic regularization Bachoc et al., (2020, 2023) but they require the use of a reference measure. Alternatively, the sliced Wasserstein distance can be used to define positive definite distance substitution kernels. It was initially proposed by Kolouri et al., (2016) in order to build kernel functions between absolutely continuous distributions with respect to the Lebesgue measure and was extented to empirical distributions by Meunier et al., (2022). Our work uses the results of Meunier et al., (2022) on the SW substitution kernel but differs in the SW approximation.
Finally, note that the notion of kernels between graphs is sometimes extended to fingerprint kernels for molecular data using for instance the Tanimoto/Jaccard kernel (Griffiths et al.,, 2024). However, such molecular kernels can only be used for this type of data and require a specific vector representation of the graphs, which is usually a complex problem.
@@ 4 SWWL GRAPH KERNEL
The proofs are given in the Supplementary Material (the positive definiteness result relies on the work of Meunier et al., (2022)). We now detail the construction of the feature map $\bm{\phi}$ involved in Eq. (5).
@@ 5.1 Classification
The proposed SWWL kernel is compared with 5 existing graph kernels, namely, WWL (Togninalli et al.,, 2019), Fused Gromov Wasserstein (FGW) (Titouan et al.,, 2019), Restricted Projected Wasserstein (RPW) (Kaloga et al.,, 2022), Propagation (PK) (Neumann et al.,, 2016) and Graph Hopper (GH) (Feragen et al.,, 2013). Five datasets are taken from the TUD benchmark (Morris et al.,, 2020): BZR, COX2, PROTEINS, ENZYMES, Cuneiform. We only keep continuous attributes for the nodes and ignore categorical information except for Cuneiform where we use the same setup as proposed by Kaloga et al., (2022) with fused attributes.
@@ 5.2 Regression
Rotor37 dataset. The NASA rotor 37 case (Roynard et al.,, 2023) serves as a prominent example of a transonic axial-flow compressor rotor widely employed in computational fluid dynamics research. This dataset is made of 3D compressible steady-state Reynold-Averaged Navier-Stokes (RANS) simulations that model external flows (Ameri,, 2009). The inputs of the simulations are given by the finite element mesh of a 3D compressor blade (graphs with $\sim 30000$ nodes), and two scalar physical parameters corresponding to the rotational speed and the input pressure. The scalar output of the problem is the isentropic efficiency, which is computed by solving the boundary value problem with the finite volume method using the elsA software (Cambier et al.,, 2011).
@@
AirfRANS dataset. This dataset consists of RANS simulations over 2D NACA airfoils with geometric variabilities in a subsonic flight regime. The inputs of the problem are given by the finite element mesh of the domain surrounding the airfoil (graphs with more than $10^{5}$ nodes), and two scalar inputs, the angle of attack and the inlet velocity. The scalar output of the problem is the drag coefficient. More details about the distributions of the training and test sets can be found in the work of Bonnet et al., (2022).
@@ Acknowledgements
This work was partially supported by the Agence Nationale de la Recherche through the SAMOURAI (Simulation Analytics and Metamodel-based solutions for Optimization, Uncertainty and Reliability AnalysIs) project under grant ANR20-CE46-0013 and the EXAMA (Methods and Algorithms at Exascale) project under grant ANR-22-EXNU-0002.
@@
Appendix A Proofs
A.1 Positive definiteness of SWWL
Definition 5 (Hilbertian (pseudo)-distance (Hein and Bousquet,, 2005)).
A pseudo-distance $d$ defined on a set $X$ is said to be Hilbertian if there exists a Hilbert space $\mathcal{H}$ and a feature map $\phi:X\rightarrow\mathcal{H}$ such that for any pair $x,x^{\prime}$ in $X$, $d(x,x^{\prime})=||\phi(x)-\phi(x^{\prime})||_{\mathcal{H}}$.
All usual kernels substituted with distances are positive definite if and only if the distance is Hilbertian as the following property shows.
Property 4 (Meunier et al., (2022)).
Introducing $\langle x,x^{\prime}\rangle^{x_{0}}_{d}=\langle\phi(x)-\phi(x_{0}),\phi(x^{\prime})-\phi(x_{0})\rangle_{\mathcal{H}}$, the following statements are equivalent:
  • $d$ is a Hilbertian pseudo-distance.
Property 5 (Peyré et al., (2019)).
$\sqrt{W_{1}}$ and $W_{2}$ are not Hilbertian distances in $\mathbb{R}^{s}$ when $s\geq 2$.
Property 6 (Meunier et al., (2022)).
$\sqrt{SW_{1}}$ and $SW_{2}$ are Hilbertian distances in $\mathbb{R}^{s}$ for all $s\geq 2$.
We can now prove the positive definiteness of SWWL.
Proof.
We recall the form of the estimated sliced Wasserstein distance between two distributions $\mu$ and $\nu$ with $P$ projections and $Q$ quantiles: $\widehat{SW}_{r,P,Q}(\mu,\nu)=||\phi(\mu)-\phi(\nu)||_{r}$ where ||.$||.||_{r}$ is the $r$-norm in $\mathbb{R}^{PQ}$ and $\bm{\phi}_{p+P(q-1)}(\mu)=(PQ)^{-1/r}F^{-1}_{\mu_{\theta_{p}}}(t_{q})$ for $p=1,\dots,P$ and $q=1,\dots,Q$. $\widehat{SW}_{r,P,Q}$ is a Hilbertian pseudo-distance 5 as we built an explicit feature map $\phi$ to $\mathcal{H}=\mathbb{R}^{PQ}$.
Finally, this property still holds if we restrict the space of measures to be $\{\mu_{G}=|V|^{-1}\sum\limits_{u\in V}\delta_{\mathbf{E}^{G}_{u}},G\in\mathcal{G}\}$, the empirical distributions associated with the continuous WL embeddings of graphs in $\mathcal{G}$.
A.2 Complexity of SWWL
In this section we give a proof of Property 2 (Complexity of SWWL).
Proof.
In comparison, using the Wasserstein distance would lead to a $O(N^{2}n^{3}\log(n))$ time complexity and a $O(n^{2})$ space complexity to build the Gram matrix, which can be prohibitive when the number of nodes is large.
Appendix B Optimal transport: definitions and properties
@@ Appendix C Details about the compared methods and their hyperparameters
In this section we give more details on hyperparameters of reference methods for classification and regression tasks presented in the experimental section. We also explain the choice of the compared methods for regression. In particular, we emphasize the difficulties of applying other approaches to the large meshes of interest, hence we also consider coarsened meshes which make it possible to compare the different methods. Note that we used indefinite kernels for the kernels methods (SVM and Gaussian process regression), even if the existence of a reproducing kernel Hilbert space in this case is not true. Some work try to extend to the indefinite case in Krein spaces, but it remains a largely open question (see for example the work of Liu et al., (2020) for kernel ridge regression). We nevertheless test several graph kernels regardless of the positive definiteness as they are commonly used and give reasonable results in practice.
@@
SWWL (ours) $\mathcal{O}(NH\delta n$ + $NPn(\log(n)+H)+N^{2}PQ)$ $P$: projections, $Q$: quantiles, $H$: WL iterations
WWL $\mathcal{O}(NH\delta n+HN^{2}n^{3}\log(n))$ $H$: WL iterations
@@
RPW $\mathcal{O}(NHn^{2}+N^{2}Pn\log(n))$ $H$: layers, $P$: last layer size
@@
Two major problems arise when comparing to WWL. The time to calculate a Wasserstein distance between embeddings of size 30000 is approximately 80 seconds, so more than 400 days are needed to compute the pairwise distances of 1000 graphs. The second issue is memory usage. Allocating matrices of size 30000 $\times$ 30000 for the transport costs of the Wasserstein distance is only possible with a minimum amount of space and severely limits parallelization. We still include a comparison with WWL on coarsened datasets with 100 parallel jobs as it is the method on which SWWL is based. Remark that we do not use entropic regularization (Cuturi, (2013)) that would reduce the complexity but has the cost of an extra hyperparameter to tune in an outer loop.
@@
The datasets Rotor37,Tensile2d, and AirfRANS can be found according to the link given in the main paper. Please note that the datasets Rotor37 that we used in the experiments and the one that is publicly available use different type of finite elements (quads/triangles) and the conversion from quads to triangles yields possibly different graphs. Scores may vary slightly with this public version.
@@ C.3 Robust Gaussian process regression
In the robust Gaussian process framework, the marginal likelihood is augmented by a prior for the range parameters $\pi(\theta,\sigma^{2},\bm{\gamma})\propto\frac{\pi(\mathbf{\gamma)}}{\sigma^{2}}$. We use the prior $\pi^{JR}(\bm{\gamma})=\left(\sum_{l=1}^{L}C_{l}\gamma_{l}^{-1}\right)^{a}\exp\left(-b\sum_{l=1}^{L}C_{l}\gamma_{l}^{-1}\right)$, $a=0.2$, $b=N^{-1/L}(a+L)$, and $C_{l}$ equal to the mean of $|x_{il}-x_{il}|$ for $1\leq i,j\leq N,i\neq j$. Remark that in the case where $x_{il}$ is not a scalar but rather a structured object, we replace the absolute value by the user-defined distance or pseudo-distance that is substituted in the kernel (in our paper we use the estimated sliced Wasserstein distance).
@@
$(\hat{\gamma_{1}},\ldots,\hat{\gamma_{L}})=\operatorname*{argmax}\limits_{\gamma_{1},\ldots,\gamma_{L}}\{\mathcal{L}(\gamma|\mathbf{X},\mathbf{y})\pi^{JR}(\bm{\gamma})\}.$ (16)
With the use of the estimated range parameters, the predictive distribution of unknown targets $\mathbf{f}_{*}:=(f(\mathbf{x}^{*}_{i}))_{i=1}^{N^{*}}$ at $N^{*}$ new locations $\mathbf{X}^{*}=(\mathbf{x}^{*}_{i})_{i=1}^{N^{*}}$ given $\mathbf{X}$, $\mathbf{y}$ and $\bm{\gamma}$ is a Student’s $t$-distribution with $N-1$ degrees of freedom
@@
with $\mathbf{h}_{*}$ being the $N^{*}\times 1$ vector composed of ones, $\hat{\theta}=(\mathbf{h}^{T}\mathbf{R}^{-1}\mathbf{h})^{-1}\mathbf{h}^{T}\mathbf{R}^{-1}\mathbf{y}$ being the least squares estimator for $\theta$, and $\mathbf{R}_{*}$, $\mathbf{R}_{**}$ being the covariance matrices evaluated between train inputs and test/train inputs, respectively.
@@
Once distance matrices are precomputed, for Rotor37 (1000 train inputs), the training takes 220 seconds; for Tensile2d (500 train inputs), the training takes 70 seconds; for AirfRANS (800 train inputs), the training takes 135 seconds. This time is similar for coarsened meshes as the shape of distance matrices is the same.
@@ D.1 Anisotropic variant of SWWL
$k_{\mathrm{ASWWL}}(G,G^{\prime})=\prod_{h=0}^{H}\exp\left(-\gamma_{h}\widehat{SW}^{2}_{2,P,Q}(\mu^{(h)}_{G},\mu^{(h)}_{G^{\prime}})\right),$ (21)